OpenStudy (anonymous):

consider the following quadratic equation: 4x^2+9x-13=-4 using the standard form ax^2+bx+c=0 of the quadratic equation above, factor the left hand side equation into two linear factors

OpenStudy (anonymous):

\[4x^2+9x-9=0\] \[(x+3)(4x-3)=0\] etc

OpenStudy (anonymous):

D=81+144=225 x1=(-9+15)/8=3/4 x2=(-9-15/8)=-3